# Dynamical tide modified Roche limit in eccentric, asynchronous binaries

### v1

@@ 3.1 Relation to the chaotic tide
We emphasize that the instability at the pericenter enters at the three-wave order (as we have demonstrated in previous sections both theoretically and numerically), a lower-order effect compared to the nonlinear anharmonicity of H. Yu et al. (2021), which fundamentally enters at the four-wave order [^4] While H. Yu et al. (2021) also considered three-wave interaction between f- and p-modes, the p-modes are not directly sourced by the tide. Therefore, to produce the anharmonicity, it requires nonlinearly sourcing the p-modes first and then letting the p-mode back-react on the f-mode; see Eq. (34). Since the three-wave coupling is used twice, it effectively becomes a four-wave interaction. The direct four-wave interaction was incorrectly ignored in H. Yu et al. (2021). There is also a numerical error in H. Yu et al. (2021) that overestimates the coupling strength between f- and p-modes.. It is also distinct from the nonlinear kick of P. Arras et al. (2023). Importantly, the correction to the tidal kick in this work is nonlinear in the fluid displacement $\boldsymbol{\xi}$, which enters the modes’ equations of motion at $\mathcal{O}(|q_{m}|^{2})\sim\mathcal{O}(M^{\prime 2}/D_{p}^{6})$; see, e.g., Eq. (20). In P. Arras et al. (2023), however, the correction is nonlinear in $(\Delta D_{p}/D_{p})$, where $\Delta D_{p}$ is the change in the pericenter separation due to tidal back-reaction. In other words, the tidal kick in P. Arras et al. (2023) is evaluated along a tidally perturbed orbit instead of the background Keplerian one. While ignored in this study, one can show that $(\Delta D_{p}/D_{p})\propto(D_{p}/M^{\prime})|q_{m}|^{2}\sim(M^{\prime}/D_{p}^{5})$ (see eq. 13 of P. Arras et al. 2023; note that the second term is smaller than the first one by $\mathcal{O}[(R_{t}/D_{p})^{3/2}]$). When evaluating $\epsilon$ at $(D_{p}+\Delta D_{p})$ instead of $D_{p}$, the correction to the tidal forcing is $\Delta\epsilon\propto(M^{\prime}/D_{p}^{3})(\Delta D_{p}/D_{p})\sim(M^{\prime 2}/D_{p}^{8})$, which is smaller than the nonlinear-in-$\boldsymbol{\xi}$ correction by $(1/D_{p}^{2})$. Since the new correction considered here is of lower order than those studied previously, the chaotic tide it leads to may occur over an expanded parameter space (especially for low eccentricities) than predicted by linear tidal kicks at the pericenter.
@@
Figure 8: Numerical evolution of the $q_{2}$ mode for a non-rotating, $n=1$ donor in an $e=0.9$ orbit with $D_{p}/R_{t}=2.9$. Different colors represent different initial amplitudes given to the mode. A large initial amplitude can be the consequence of the dynamical tide’s chaotic, long-term evolution, and it allows runaways in mode amplitude to occur at $D_{p}/R_{t}$ greater than Fig. 6.
@@
In turn, for highly eccentric systems, the chaotic tide can make the threshold pericenter separation greater than the prediction of Fig. 6. Note that in Fig. 6, the modes are initiated by their leading-order equilibrium tide solution at the apocenter. For a fixed $D_{p}$, a large eccentricity essentially suppresses the initial amplitude by a factor of $(1+e)/(1-e)]^{3}$. Over merely 22 orbits, the chaotic growth is limited as, on average, the mode energy increases linearly with the number of passages (Y. Wu, 2018; M. Vick & D. Lai, 2018). One can nonetheless capture the long-term growth history of the modes by giving them different initial conditions. One such example is shown in Fig. 8. We consider in this example the non-rotating $n=1$ model with $e=0.9$ and $D_{p}/R_{t}=2.9$. Starting with the equilibrium amplitude, the mode amplitude remains oscillatory within 22 orbits (the bluest curve; also Fig. 6). However, if a significant initial amplitude is used (indicated by curves with different colors), the mode may run away at this large separation. As the chaotic growth is almost unbounded (until reaching near equipartition with the orbit), the threshold $D_{p}/R_{t}$ for the mode’s runaway can be as large as that for the chaotic tide to occur.
@@
In fact, the runaways in our simulation correspond to the catastrophic wave breakings during the chaotic evolution postulated in Y. Wu (2018). When a mode’s amplitude runs away, it cannot be supported by the background and has to break somehow. However, our model does not describe the consequence of such a break. One possibility, as suggested in M. MacLeod et al. (2022), would be shock-heating the surface layer with some potential mass ejection from the donor. It further suggests that energy dissipation in highly eccentric systems can be very efficient due to not a continuous dissipation as in the linear theory but discrete epochs of chaotic growth followed by wave breakings as pointed out by Y. Wu (2018). Meanwhile, mass loss is likely nonconservative based on J. Guillochon et al. (2011). More detailed analysis (theoretical or numerical) of the saturation of the instability is left for future investigations.
3.2 High eccentricity migration of exoplanets
@@
Related to specific astrophysical settings, the first application we consider is the formation of HJs via the high-eccentricity migration channel (see, e.g., R. I. Dawson & J. A. Johnson 2018 for a review). In particular, our model provides a theoretical explanation for why J. Guillochon et al. (2011) observed mass loss from a migrating Jovian planet at a pericenter separation as large as $D_{p}\simeq 2.7R_{t}$, $35\%$ greater than the Roche-lobe-overfilling separation of $D_{p}\simeq 2.0R_{t}$ using the fit by P. P. Eggleton (1983). Extrapolating the low-eccentricity Roche lobe calculation of J. F. Sepinsky et al. (2007) will increase the discrepancy even further. We highlight that our Fig. 5 shows remarkable similarities to the simulations of J. Guillochon et al. (2011); see, e.g., their fig. 3. Specifically, there should be no significant mass loss from $M$ (the planet) in a single passage. Instead, the perturbation accumulates over orbits, and significant mass loss from $M$ eventually occurs after $\mathcal{O}(10)$ orbits. This consistency validates our analysis.
@@
More importantly, our theoretical model (in terms of simple ordinary differential equations; Eq. 19 or 15) allows the more expensive three-dimensional hydrodynamical simulations of J. Guillochon et al. (2011) to be faithfully extrapolated to different systems based on each system’s unique configuration. In particular, the commonly adopted disruption threshold of $2.7R_{t}$ (e.g., K. R. Anderson et al. 2016; M. Vick et al. 2019) is only true for the specific system studied in J. Guillochon et al. (2011) over $\mathcal{O}(10)$ passages. More orbits of integration may increase the threshold further due to the chaotic build-up of the mode (captured by a non-equilibrium initial condition shown in Fig. 8). More interestingly, if either the orbital eccentricity or the planet’s rotational rate varies, a different disruption threshold would also apply; see Figs. 6 and 4. The dependence of $D_{p}^{\rm(th)}$ on these properties may indicate interesting and unanticipated correlations between the current HJ distribution and progenitor configurations.
@@
Our Fig. 8 further suggests that the Roche limit for highly eccentric systems can be as large as the threshold separation to trigger the chaotic tide at about $3.8R_{t}$ (Y. Wu, 2018) for a typical proto-HJ. Interestingly, this distance coincides with the location where a pile-up of HJs is observed (with a current period of about 3 days; Y. Wu 2018; R. I. Dawson & J. A. Johnson 2018). On the other hand, HJ systems with separations shorter than this could imply a high spin rate of the proto-HJ (young system), as the thresholds for both chaotic tide (e.g., fig. 4 of H. Yu et al. 2021) and Roche limit (Fig. 4) are suppressed by rotation.
@@
Our Fig. 6 may also have implications for formation channels other than the standard high-eccentricity migration. For example, a planet may have disk-migrated before experiencing a high-eccentricity migration. In such a hybrid scenario, the eccentricity excitation required in the second stage is only moderate, with $e<0.9$ (e.g., figs. 5 and 10 of H. Yu & F. Dai 2024). This two-stage migration is suggested by J. Guillochon et al. (2011) based on observations available at the time and supported by the recent investigation of H. Yu & F. Dai (2024) on the formation of the WASP-107 system (their section 4.3). Both studies found that the initial apocenter distance of the planet of the high-eccentricity migration stage is within the snow line. Our Fig. 6 suggests that the HJ/HN formation efficiency for such a system with $e<0.9$ may be higher than originally expected based on a fixed $D_{p}^{(\rm th)}=2.7R_{t}$ derived from $e=0.9$. An increase in the predicted HJ formation rate will improve the agreement between theory and observation (see, e.g., table 2 of R. I. Dawson & J. A. Johnson 2018).
@@
As discussed above in §3.1, the nonlinear tidal interactions leading to the Roche limit can also help trigger and maintain the chaotic tide. For highly eccentric systems with $e>0.9$, an expanded parameter space for the chaotic tide is crucial for explaining the lack of super-eccentric proto-HJs A. Socrates et al. (2012); R. I. Dawson et al. (2015). The efficient energy dissipation due to the chaotic tide naturally explains the paucity, yet the original study of Y. Wu (2018) applies only to those with current periods less than 3 days, a small fraction of the missing population (R. I. Dawson & J. A. Johnson, 2018). Expanding the threshold for the chaotic tide is therefore crucial for reducing the tension with observations. For those with $e<0.9$, the propagation phase of a mode itself is not sufficiently random, and the nonlinear instability at the pericenter is likely what maintains the chaotic tide and drives the orbital evolution. It is important to note that there is no consensus on the dissipation mechanism in moderately eccentric systems after the quench of the chaotic tide. The predicted dissipation rate of f-modes from first-principle calculations (P. Goldreich & P. D. Nicholson, 1977) is orders of magnitude below what is needed for a planet to circularize without the chaotic tide (but see, e.g., Y. Wu 2005; G. I. Ogilvie 2014).
@@
Repeating X-ray bursts at galactic centers provide valuable information about the properties of the central MBH and its surrounding environments. A rpTDE corresponds to a star partially disrupted by the tidal force of the MBH, with the remnant continuing to orbit and getting disrupted again in the future. Compared to a full TDE, rpTDEs have a larger $D_{p}/R_{t}$, and a different mass fallback rate E. R. Coughlin & C. J. Nixon (2019). Through a hydrodynamic simulation, T. Ryu et al. (2020) finds that a main sequence star undergoes a partially disrupting instead of a full TDE for $2\gtrsim D_{p}/R_{t}\gtrsim 0.5$ for a single passage in the Newtonian limit. S. Rosswog et al. (2009a) performs a simulation of the tidal disruption of a WD orbiting a $10^{3}~M_{\odot}$ intermediate mass black hole, where mass loss occurs at $1.1R_{t}$ in a single passage. Unlike full TDEs that are typically assumed to happen on nearly parabolic orbits, rpTDEs are likely to have more moderate eccentricities with $1-e\sim 0.01$ (K. Hayasaki et al., 2013; M. Cufari et al., 2022; Z. Liu et al., 2024). The typical orbital period is not long enough to damp out the dynamical tide (unless the quality factor is significantly lower than $10^{6}$; see the discussions at the end of §2.3), allowing it to build up over multiple orbits and potentially change the single-passage predication significantly. Similar to the case with Jovian planets, our model provides a theoretical model to describe the disruption, accounting for the accumulation history of the dynamical tide over multiple orbits. As the tidal amplitude builds up chaotically until it runs away and ejects mass to power a burst, it can be a promising mechanism to produce rpTDEs with irregular recurrence time (S. Y. Lau & H. Yu, 2025), which has been observed in, e.g., J0456-20 (Z. Liu et al., 2024) and HLX-1 (O. Godet et al., 2014; N. A. Webb et al., 2023). Moreover, the threshold pericenter separation $D_{p}^{(\rm th)}$ for the disruption to marginally occur may increase by $\gtrsim 60\%$ with the buildup (comparing Figs. 6 and 7 at $e=0.9$). Since $D_{p}^{\rm(th)}$ sets the loss cone radius, an increase in $D_{p}^{(\rm th)}$ increases the rpTDE event rate (E. Bortolas et al., 2023). If we assume an isotropic distribution of a star’s velocity vector at a fixed distance for simplicity, the increase in loss cone event rates is proportional to the increase in its radius (D. Merritt, 2013).
@@
Our results also complement some previous models on QPEs. In A. King (2022); M. Wang et al. (2022), it is proposed that a highly eccentric WD-MBH binary can simultaneously produce TDEs and QPEs. The predicted tidal stripping based on Eggleton’s formula is strictly periodic with $D_{p}<2R_{t}$. The QPEs produced via this mechanism would therefore be highly regular, which contradicts the observed fluctuations in the recurrence time of bursts. The stochastic nature of the mode provides a natural explanation for these fluctuations. The chaotic evolution also means that the tidal disruption does not repeat every orbit, as the mean mode energy builds up diffusively over time. Its corresponding impact on the orbital evolution and mass transfer has been studied in S. Y. Lau & H. Yu (2025). A significant shortcoming within this work is that although the parametrized mass loss formalism is flexible enough to model the case with dynamical tide, the range of the parameters is prescribed to match the observations and previous hydrodynamic simulation results in J. Guillochon et al. (2011). In other words, it still lacks a direct connection to the first-principles calculation considering the hydrodynamics under an external tidal field. This paper provides a theoretical description of the development from mode oscillations into a runaway process that can potentially lead to mass loss.
@@
For a deep encounter ($D_{p}/R_{t}<0.5$) of a WD with the MBH, a thermonuclear runaway can be triggered by the tidal compression J. P. Luminet & B. Pichon (1989); S. Rosswog et al. (2009b); M. MacLeod et al. (2016); A. Tanikawa (2018), depending on its composition. Based on the model by B. Carter & J. P. Luminet (1983), the star undergoes a phase in which it becomes flattened in the orbital plane and the interior is strongly compressed. J. P. Luminet & B. Pichon (1989) shows that it can trigger nucleosynthesis processes similar to that at the core of a main-sequence star. The energy release can be comparable to that of Type 1a supernovae S. Rosswog et al. (2009b). This scenario is favored for lower mass MBHs, of $\sim 10^{3}M_{\odot}$, where the WD orbit can penetrate deeper without getting directly swallowed. Whether a WD ignites or loses mass and powers rpTDE depends on how $\Delta E_{\rm mode}$ from a runaway is released. If $\Delta E_{\rm mode}$ is mostly used to eject mass, then rpTDE is the outcome. If, however, it is mostly converted into heat in the core, then the WD may be ignited in a thermonuclear explosion. Determining the outcome is beyond the scope of the current model, but an interesting question to follow up on in the future.
4 Conclusion and discussion

### v2

@@ III.1 Relation to the diffusive tide
We emphasize that the instability at the pericenter enters at the three-wave order (as we have demonstrated in previous sections both theoretically and numerically), a lower-order effect compared to the nonlinear anharmonicity of Yu et al. (2021), which fundamentally enters at the four-wave order [^5] While Yu et al. (2021) also considered three-wave interaction between f- and p-modes, the p-modes are not directly sourced by the tide. Therefore, to produce the anharmonicity, it requires nonlinearly sourcing the p-modes first and then letting the p-mode back-react on the f-mode; see Eq. (36). Since the three-wave coupling is used twice, it effectively becomes a four-wave interaction. The direct four-wave interaction was incorrectly ignored in Yu et al. (2021). There is also a numerical error in Yu et al. (2021) that overestimates the coupling strength between f- and p-modes.. It is also distinct from the nonlinear kick of Arras et al. (2023). Importantly, the correction to the tidal kick in this work is nonlinear in the fluid displacement $\boldsymbol{\xi}$, which enters the modes’ equations of motion at $\mathcal{O}(|q_{m}|^{2})\sim\mathcal{O}(M^{\prime 2}/D_{p}^{6})$; see, e.g., Eq. (21). In Arras et al. (2023), however, the correction is nonlinear in $(\Delta D_{p}/D_{p})$, where $\Delta D_{p}$ is the change in the pericenter separation due to tidal back-reaction. In other words, the tidal kick in Arras et al. (2023) is evaluated along a tidally perturbed orbit instead of the background Keplerian one. While ignored in this study, one can show that $(\Delta D_{p}/D_{p})\propto(D_{p}/M^{\prime})|q_{m}|^{2}\sim(M^{\prime}/D_{p}^{5})$ (see eq. 13 of Arras et al. 2023; note that the second term is smaller than the first one by $\mathcal{O}[(R_{t}/D_{p})^{3/2}]$). When evaluating $\epsilon$ at $(D_{p}+\Delta D_{p})$ instead of $D_{p}$, the correction to the tidal forcing is $\Delta\epsilon\propto(M^{\prime}/D_{p}^{3})(\Delta D_{p}/D_{p})\sim(M^{\prime 2}/D_{p}^{8})$, which is smaller than the nonlinear-in-$\boldsymbol{\xi}$ correction by $(1/D_{p}^{2})$. Since the new correction considered here is of lower order than those studied previously, the diffusive (chaotic) tide it leads to may occur over an expanded parameter space (especially for low eccentricities) than predicted by linear tidal kicks at the pericenter.
@@
Figure 8: Numerical evolution of the $q_{2}$ mode for a non-rotating, $n=1$ donor in an $e=0.9$ orbit with $D_{p}/R_{t}=2.9$. Different colors represent different initial amplitudes given to the mode. A large initial amplitude can be the consequence of the dynamical tide’s diffusive, long-term evolution, and it allows runaways in mode amplitude to occur at $D_{p}/R_{t}$ greater than Fig. 6.
@@
In turn, for highly eccentric systems, the diffusive tide can make the threshold pericenter separation greater than the prediction of Fig. 6. Note that in Fig. 6, the modes are initiated by their leading-order equilibrium tide solution at the apocenter. For a fixed $D_{p}$, a large eccentricity essentially suppresses the initial amplitude by a factor of $(1+e)/(1-e)]^{3}$. Over merely 22 orbits, the diffusive growth is limited as, on average, the mode energy increases linearly with the number of passages (Wu, 2018; Vick and Lai, 2018). One can nonetheless capture the long-term growth history of the modes by giving them different initial conditions. One such example is shown in Fig. 8. We consider in this example the non-rotating $n=1$ model with $e=0.9$ and $D_{p}/R_{t}=2.9$. Starting with the equilibrium amplitude, the mode amplitude remains oscillatory within 22 orbits (the bluest curve; also Fig. 6). However, if a significant initial amplitude is used (indicated by curves with different colors), the mode may run away at this large separation. As the diffusive growth is almost unbounded (until reaching near equipartition with the orbit), the threshold $D_{p}/R_{t}$ for the mode’s runaway can be as large as that for the diffusive tide to occur.
@@
In fact, the runaways in our simulation correspond to the catastrophic wave breakings during the diffusive evolution postulated in Wu (2018). When a mode’s amplitude runs away, it cannot be supported by the background and has to break somehow. However, our model does not describe the consequence of such a break. One possibility, as suggested in MacLeod et al. (2022), would be shock-heating the surface layer with some potential mass ejection from the donor. It further suggests that energy dissipation in highly eccentric systems can be very efficient due to not a continuous dissipation as in the linear theory but discrete epochs of diffusive growth followed by wave breakings as pointed out by Wu (2018). Meanwhile, mass loss is likely nonconservative based on Guillochon et al. (2011). More detailed analysis (theoretical or numerical) of the saturation of the instability is left for future investigations.
III.2 High eccentricity migration of exoplanets
@@
Related to specific astrophysical settings, the first application we consider is the formation of HJs via the high-eccentricity migration channel (see, e.g., Dawson and Johnson 2018 for a review). In particular, our model provides a theoretical explanation for why Guillochon et al. (2011) observed mass loss from a migrating Jovian planet at a pericenter separation as large as $D_{p}\simeq 2.7R_{t}$, $35\%$ greater than the Roche-lobe-overfilling separation of $D_{p}\simeq 2.0R_{t}$ using the fit by Eggleton (1983). Extrapolating the low-eccentricity Roche lobe calculation of Sepinsky et al. (2007) will increase the discrepancy even further. We highlight that our Fig. 5 shows remarkable similarities to the simulations of Guillochon et al. (2011); see, e.g., their fig. 3. Specifically, there should be no significant mass loss from $M$ (the planet) in a single passage. Instead, the perturbation accumulates over orbits, and significant mass loss from $M$ eventually occurs after $\mathcal{O}(10)$ orbits. This consistency validates our analysis.
@@
More importantly, our theoretical model (in terms of simple ordinary differential equations; Eq. 20 or 16) allows the more expensive three-dimensional hydrodynamical simulations of Guillochon et al. (2011) to be faithfully extrapolated to different systems based on each system’s unique configuration. In particular, the commonly adopted disruption threshold of $2.7R_{t}$ (e.g., Anderson et al. 2016; Vick et al. 2019) is only true for the specific system studied in Guillochon et al. (2011) over $\mathcal{O}(10)$ passages. More orbits of integration may increase the threshold further due to the diffusive build-up of the mode (captured by a non-equilibrium initial condition shown in Fig. 8). Recent studies by Yu and Dai (2024); Weldon et al. (2025) adopted a (potentially more realistic) soft disruption prescription where the planetary mass loss per orbit is parameterized as an exponentially decaying function of $D_{p}/R_{t}$, yet the parameters are still calibrated to Guillochon et al. (2011) and held constant throughout the evolution. If either the orbital eccentricity or the planet’s rotational rate varies, a different disruption threshold (or a different set of mass loss prescriptions) should apply; see Figs. 6 and 4. The dependence of $D_{p}^{\rm(th)}$ on these properties may indicate interesting and unanticipated correlations between the current HJ distribution and progenitor configurations.
@@
Our Fig. 8 further suggests that the Roche limit for highly eccentric systems can be as large as the threshold separation to trigger the diffusive tide at about $3.8R_{t}$ (Wu, 2018) for a typical proto-HJ. Interestingly, this distance coincides with the location where a pile-up of HJs is observed (with a current period of about 3 days; Wu 2018; Dawson and Johnson 2018). On the other hand, HJ systems with separations shorter than this could imply a high spin rate of the proto-HJ (young system), as the thresholds for both diffusive tide (e.g., fig. 4 of Yu et al. 2021) and Roche limit (Fig. 4) are suppressed by rotation.
@@
Our Fig. 6 may also have implications for formation channels other than the standard high-eccentricity migration. For example, a planet may have disk-migrated before experiencing a high-eccentricity migration. In such a hybrid scenario, the eccentricity excitation required in the second stage is only moderate, with $e<0.9$ (e.g., figs. 5 and 10 of Yu and Dai 2024). This two-stage migration is suggested by Guillochon et al. (2011) based on observations available at the time and supported by the recent investigation of Yu and Dai (2024) on the formation of the WASP-107 system (their section 4.3). Both studies found that the initial apocenter distance of the planet of the high-eccentricity migration stage is within the snow line. Our Fig. 6 suggests that the HJ/HN formation efficiency for such a system with $e<0.9$ may be higher than originally expected based on a fixed $D_{p}^{(\rm th)}=2.7R_{t}$ derived from $e=0.9$. An increase in the predicted HJ formation rate will improve the agreement between theory and observation (see, e.g., table 2 of Dawson and Johnson 2018).
@@
As discussed above in §III.1, the nonlinear tidal interactions leading to the Roche limit can also help trigger and maintain the diffusive tide. For highly eccentric systems with $e>0.9$, an expanded parameter space for the diffusive tide is crucial for explaining the lack of super-eccentric proto-HJs Socrates et al. (2012); Dawson et al. (2015). The efficient energy dissipation due to the diffusive tide naturally explains the paucity, yet the original study of Wu (2018) applies only to those with current periods less than 3 days, a small fraction of the missing population (Dawson and Johnson, 2018). Expanding the threshold for the diffusive tide is therefore crucial for reducing the tension with observations. For those with $e<0.9$, the propagation phase of a mode itself is not sufficiently random, and the nonlinear instability at the pericenter is likely what maintains the diffusive tide and drives the orbital evolution. It is important to note that there is no consensus on the dissipation mechanism in moderately eccentric systems after the quench of the diffusive tide. The predicted dissipation rate of f-modes from first-principle calculations (Goldreich and Nicholson, 1977) is orders of magnitude below what is needed for a planet to circularize without the diffusive tide (but see, e.g., Wu 2005; Ogilvie 2014).
@@
Repeating X-ray bursts at galactic centers provide valuable information about the properties of the central MBH and its surrounding environments. A rpTDE corresponds to a star partially disrupted by the tidal force of the MBH, with the remnant continuing to orbit and getting disrupted again in the future. Compared to a full TDE, rpTDEs have a larger $D_{p}/R_{t}$, and a different mass fallback rate Coughlin and Nixon (2019). Through a hydrodynamic simulation, Ryu et al. (2020) finds that a main sequence star undergoes a partially disrupting instead of a full TDE for $2\gtrsim D_{p}/R_{t}\gtrsim 0.5$ for a single passage in the Newtonian limit. Rosswog et al. (2009b) performs a simulation of the tidal disruption of a WD orbiting a $10^{3}~M_{\odot}$ intermediate mass black hole, where mass loss occurs at $1.1R_{t}$ in a single passage. Unlike full TDEs that are typically assumed to happen on nearly parabolic orbits, rpTDEs are likely to have more moderate but still significant eccentricities with $1-e\sim 0.1-0.01$ (Hayasaki et al., 2013; Guillochon and Ramirez-Ruiz, 2013; Cufari et al., 2022; Liu et al., 2024; Bandopadhyay et al., 2024, 2025). The long orbital period limits previous numerical simulations to binaries with $D_{p}/R_{t}\lesssim 2$ in order for mass loss to occur in a single pericenter passage. However, the orbital period is not long enough to damp out the dynamical tide (unless the quality factor is significantly lower than $10^{6}$ in which case the study of Yao and Quataert 2025 applies; see the discussions at the end of §II.3), allowing it to build up over multiple orbits and potentially change the single-passage predication significantly. Similar to the case with Jovian planets, our study provides a new, theoretical model to describe the disruption, accounting for the accumulation history of the dynamical tide over multiple orbits. In our picture, the dynamical tide builds up diffusively until it runs away and ejects mass to power a burst. The stochastic growth can be a promising mechanism to produce rpTDEs with irregular recurrence times (Lau and Yu, 2025, 2026), which has been observed in, e.g., J0456-20 (Liu et al., 2024) and HLX-1 (Godet et al., 2014; Webb et al., 2023) that previous models requiring mass loss at every pericenter passage struggle to reproduce. For example, the simulations in Liu et al. (2024) require a fractional mass loss of almost 90% to reproduce the $\sim 30\%$ fluctuations in the recurrence times of J0456-20 from $\sim 300\,{\rm days}$ to $\sim 200\,{\rm days}$. Furthermore, this amount of mass loss is also in tension with the inferred eruption luminosity.
Moreover, the threshold pericenter separation $D_{p}^{(\rm th)}$ for the disruption to marginally occur may increase by $\gtrsim 60\%$ with the buildup compared to the single-passage predictions (e.g., Guillochon and Ramirez-Ruiz 2013; Ryu et al. 2020; comparing Figs. 6 and 7 at $e=0.9$). Since $D_{p}^{\rm(th)}$ sets the loss cone radius, an increase in $D_{p}^{(\rm th)}$ increases the rpTDE event rate (Bortolas et al., 2023). If we assume an isotropic distribution of a star’s velocity vector at a fixed distance for simplicity, the increase in loss cone event rates is proportional to the increase in its radius (Merritt, 2013).
@@
Our results also complement some previous models on QPEs. In King (2022); Wang et al. (2022), it is proposed that a highly eccentric WD-MBH binary can simultaneously produce TDEs and QPEs. The predicted tidal stripping based on Eggleton’s formula is strictly periodic with $D_{p}<2R_{t}$. The QPEs produced via this mechanism would therefore be highly regular, which contradicts the observed fluctuations in the recurrence time of bursts. The stochastic nature of the mode provides a natural explanation for these fluctuations. The diffusive evolution also means that the tidal disruption does not repeat every orbit, as the mean mode energy builds up diffusively over time. Its corresponding impact on the orbital evolution and mass transfer has been studied in Lau and Yu (2025). A significant shortcoming within this work is that although the parametrized mass loss formalism is flexible enough to model the case with dynamical tide, the range of the parameters is prescribed to match the observations and previous hydrodynamic simulation results in Guillochon et al. (2011). In other words, it still lacks a direct connection to the first-principles calculation considering the hydrodynamics under an external tidal field. This paper provides a theoretical description of the development from mode oscillations into a runaway process that can potentially lead to mass loss.
@@
For a deep encounter ($D_{p}/R_{t}<0.5$) of a WD with the MBH, a thermonuclear runaway can be triggered by the tidal compression Luminet and Pichon (1989); Rosswog et al. (2009a); MacLeod et al. (2016); Tanikawa (2018), depending on its composition. Based on the model by Carter and Luminet (1983), the star undergoes a phase in which it becomes flattened in the orbital plane and the interior is strongly compressed. Luminet and Pichon (1989) shows that it can trigger nucleosynthesis processes similar to that at the core of a main-sequence star. The energy release can be comparable to that of Type 1a supernovae Rosswog et al. (2009a). This scenario is favored for lower mass MBHs, of $\sim 10^{3}M_{\odot}$, where the WD orbit can penetrate deeper without getting directly swallowed. Whether a WD ignites or loses mass and powers rpTDE depends on how $\Delta E_{\rm mode}$ from a runaway is released. If $\Delta E_{\rm mode}$ is mostly used to eject mass, then rpTDE is the outcome. If, however, it is mostly converted into heat in the core, then the WD may be ignited in a thermonuclear explosion. Determining the outcome is beyond the scope of the current model, but an interesting question to follow up on in the future.
IV Conclusion and discussion